Triangle ABC with vertices A(2,3), B(0,4), C(-3,-3) is Translated 3 units right and 1 unit down.
Write Translation matrix.
Find corrdinates.
How do you solve this?
Thanks.

Answer by checkley71(8403)

A(2,3) MOVED 3 TO THE RIGHT & DOWN 1 IS:
A(2+3,3-1)=(5,2)
B(0,4) BECOMES (0+3,4-1)=(3,3)
C(-3,-3) BECOMES (-3+3,-3-1)=(0,-4)
TO PROVE THIS USE GRAPH PAPER & PLOT ALL THESE POINTS.